# LoRA-Pro: Are Low-Rank Adapters Properly Optimized?

### v1

@@ Appendix B Proof of Theoretical Results
Theorem B.2 When updating matrices $A$ and $B$ using the closed-form solution from Theorem 3.3, we proceed as follows: $A\leftarrow A-\gamma g^{A},$ (26) $B\leftarrow B-\gamma g^{B},$ (27) where $\gamma\geq 0$ denotes the learning rate. Our method ensures a decrease in the loss, akin to the standard gradient descent algorithm, expressed by: $\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}\leq 0$ (28)
Proof (Part 1) In summary, the proof of Theorem 3.3 is divided into two distinct parts. To begin with, we demonstrate that $\mathrm{d}L$ can be expressed in the following form: $\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}.$ (29) In the second part, we prove that this expression for $\mathrm{d}L$ is always less than or equal to zero: $\mathrm{d}L\leq 0$. Therefore, in this part, we first prove Equation (29). During the optimization process, the differential change in the loss function, $\mathrm{d}L$, can be expressed in terms of the differentials $\mathrm{d}A$ and $\mathrm{d}B$ as follows: $\mathrm{d}L=\langle\frac{\partial L}{\partial A},\mathrm{d}A\rangle_{F}+\langle\frac{\partial L}{\partial B},\mathrm{d}B\rangle_{F}.$ (30) From Equation (26) and (27), we can derive that: $\mathrm{d}A=-\gamma g^{A},\quad\mathrm{d}B=\gamma g^{B}.$ (31) Given that $\frac{\partial L}{\partial A}=g^{A}_{lora}$ and $\frac{\partial L}{\partial B}=g^{B}_{lora}$, it follows that: $\mathrm{d}L$ $=-\gamma(\langle g^{A}_{lora},g^{A}\rangle_{F}+\langle g^{B}_{lora},g^{B}\rangle_{F})$ (32) $=-\gamma(\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}$ $+\langle g^{A}_{lora},XA\rangle_{F}-\langle g^{B}_{lora},BX\rangle_{F}).$ And we have the following equation: $\langle g^{A}_{lora},XA\rangle_{F}-\langle g^{B}_{lora},BX\rangle_{F}$ (33) $=$ $\langle g^{A}_{lora}A^{T},X\rangle_{F}-\langle B^{T}g^{B}_{lora},X\rangle_{F}$ $=$ $\langle g^{A}_{lora}A^{T}-B^{T}g^{B}_{lora},X\rangle_{F}$ $=$ $\langle(sB^{T}g)A^{T}-B^{T}(sgA^{T}),X\rangle_{F}$ $=$ $0$ Therefore, we have: $\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}.$ (34)
Proof (Part 2) In this part, we aim to prove $\mathrm{d}L\leq 0$. Given that the learning rate $\gamma>0$, it suffices to show the following inequalities: $\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}\geq 0,$ (35) $\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\geq 0.$ (36) By proving these inequalities, we can establish that $\mathrm{d}L\leq 0$ as derived from Equation (29). ① Proof of $\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}\geq 0$. To begin with, we need to show that $(B^{T}B)^{-1}$ is positive definite. To establish this, it is sufficient to show that $B^{T}B$ is positive definite, as the inverse of a positive definite matrix is also positive definite. To achieve this, consider any non-zero vector $x$, and noting that $B$ is full-rank, we have, $\langle x,B^{T}Bx\rangle=\langle Bx,Bx\rangle=\|Bx\|^{2}>0.$ (37) This shows that $B^{T}B$ is positive definite. Consequently, $(B^{T}B)^{-1}$ is positive definite as well. Since $(B^{T}B)^{-1}$ is positive definite, and thus we can apply Cholesky decomposition, and $(B^{T}B)^{-1}=UU^{T}$. With this, we have, $\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}$ $=\frac{1}{s^{2}}\langle g^{A}_{lora},UU^{T}g^{A}_{lora}\rangle_{F}$ (38) $=\frac{1}{s^{2}}\langle U^{T}g^{A}_{lora},U^{T}g^{A}_{lora}\rangle_{F}$ $=\frac{1}{s^{2}}\|U^{T}g^{A}_{lora}\|_{F}^{2}\geq 0$ ② Proof of $\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\geq 0$. Similarly, we can prove that matrix $(AA^{T})^{-1}$ is positive-definite. By employing Cholesky decomposition, we express $(AA^{T})^{-1}=UU^{T}$, where $U$ is a lower-triangle matrix. Subsequently, we define $P=I-B(B^{T}B)^{-1}B^{T}$. It can be shown that $P^{2}=P$, indicating that $P$ is a projection matrix. Consequently, the eigenvalues of $P$ are either 0 or 1, which implies that $P$ is positive semi-definite. Utilizing the Cholesky decomposition, we derive that $P=VV^{T}$, where $V$ is a lower-triangle matrix. Finally, we have: $\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}$ $=\frac{1}{s^{2}}\langle g^{B}_{lora},VV^{T}g^{B}_{lora}UU^{t}\rangle_{F}$ (39) $=\frac{1}{s^{2}}\langle V^{T}g^{B}_{lora}U,V^{T}g^{B}_{lora}U\rangle_{F}$ $=\frac{1}{s^{2}}\|V^{T}g^{B}_{lora}U\|_{F}^{2}\geq 0$ In summary, based on the above proofs, we have demonstrated that: $\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}\leq 0$ (40)
@@
Theorem B.3 Consider the optimization problem, $\min_{X}\|g^{A}-g^{A}_{lora}\|_{F}^{2}+\|g^{B}-g^{B}_{lora}\|_{F}^{2},$ (41) where $g^{A}$ and $g^{B}$ are the optimal solutions as stated in Theorem 3.3. The optimal $X$ can be determined by solving the Sylvester equation: $B^{T}BX+XAA^{T}=-\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}A^{T},$ (42) which has a unique solution $X$ provided that $B^{T}B$ and $-AA^{T}$ do not have any shared eigenvalues.
Proof For simplicity, we denote $L=\|g^{A}-g^{A}_{lora}\|_{F}^{2}+\|g^{B}-g^{B}_{lora}\|_{F}^{2}$. To solve the optimization problem, we need to satisfy the following conditions: $\frac{\partial L}{\partial X}=0.$ (43) Since $g^{A}$ and $g^{B}$ are solutions in Theorem 3.3 and $g^{A}_{lora}=sB^{T}g$ and $g^{B}_{lora}=sgA^{T}$, we obtain that: $2(g^{A}-g^{A}_{lora})A^{T}$ $-2B^{T}(g^{B}-g^{B}_{lora})=0,$ (44) $\Rightarrow\quad g^{A}A^{T}-B^{T}g^{B}$ $=g^{A}_{lora}A^{T}-B^{T}g^{B}_{lora},$ $\Rightarrow\quad B^{T}BX+XAA^{T}$ $=-\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}A^{T},$ which is a Sylvester equation. This equation has a unique solution for $X$ if and only if $B^{T}B$ and $-AA^{T}$ have no shared eigenvalues.
@@ Appendix C Optimization Algorithms
In this section, we present the pseudo-codes for implementing our LoRA-Pro method using the SGD [Sutskever et al., 2013] and AdamW [Loshchilov and Hutter, 2019] optimizers. These are detailed in Algorithm 1 and Algorithm 2, respectively.
In the standard SGD algorithm, as illustrated in Algorithm 1, all we need to do is adjusting the gradients of matrices $A$ and $B$ with the solutions in Theorem 3.3.
In AdamW optimizer, the implementation becomes more complex. Several modifications are necessary. Firstly, in order to mimic full fine-tuning, after adjusting the gradients of matrices $A$ and $B$, we need to compute the equivalent gradient,
$\tilde{g}=sg^{B}A+sBg^{A}.$ (45)
$W\leftarrow(1-\gamma\lambda)(W_{0}+sBA).$ (47)
This can be decomposed into:
@@
Algorithm 2 LoRA-Pro with AdamW optimizer

### v2

@@ B.1 Proof that the Equivalent Gradient is Low-Rank
Theorem.
Assume matrices $B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}$ are both full rank. For the objective $\min_{g^{A},g^{B}}\|\tilde{g}-g\|^{2}_{F}$, the solutions are given by: $g^{A}$ $=\frac{1}{s}(B^{T}B)^{-1}B^{T}g+XA=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}+XA$ (24) $g^{B}$ $=\frac{1}{s}[I-B(B^{T}B)^{-1}B^{T}]gA^{T}(AA^{T})^{-1}-BX$ (25) $=\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}-BX.$ (26) Here, $X\in\mathbb{R}^{r\times r}$ represents an arbitrary matrix.
Proof.
For simplicity, we denote $L=\|sBg^{A}+sg^{B}A-g\|_{F}^{2}$. To solve the optimization problem, we need to satisfy the following conditions:
$\frac{\partial L}{\partial g^{A}}$ $=2sB^{T}(sBg^{A}+sg^{B}A-g)=0$ (27)
$\frac{\partial L}{\partial g^{B}}$ $=2(sBg^{A}+sg^{B}A-g)sA^{T}=0$ (28)
Given that matrices $A$ and $B$ are full-rank, $AA^{T}$ and $B^{T}B$ are invertible. And from Equation (28), we derive:
$g^{B}=\frac{1}{s}gA^{T}(AA^{T})^{-1}-Bg^{A}A^{T}(AA^{T})^{-1}.$ (29)
Substituting this into Equation (27), we obtain the following linear equation:
$g^{A}[I-A^{T}(AA^{T})^{-1}A]=\frac{1}{s}(B^{T}B)^{-1}B^{T}g[I-A^{T}(AA^{T})^{-1}A].$ (30)
Here, we notice that the matrix $P=I-A^{T}(AA^{T})^{-1}A$ is a projection matrix with rank $n-r$. The solution to the linear equation (30) is:
$g^{A}=\frac{1}{s}(B^{T}B)^{-1}B^{T}g+XA,$ (31)
where $X\in\mathbb{R}^{r\times r}$ represents an arbitrary matrix. We take the solution (31) into Equation (29), we derive:
$g^{B}=\frac{1}{s}[I-B(B^{T}B)^{-1}B^{T}]gA^{T}(AA^{T})^{-1}-BX$ (32)
While we have obtained closed-form solutions for $g^{A}$ and $g^{B}$, these solutions explicitly depend on the gradient of the matrix $W$, i.e., $g$, which is undesirable since $g$ is unknown during LoRA optimization. Fortunately, the solutions can be transformed into the forms of the gradients of standard LoRA, where the gradients are:
$g^{A}_{lora}=sB^{T}g,\quad g^{B}_{lora}=sgA^{T}.$ (33)
Therefore, the solutions to the optimization problem can be written as:
$g^{A}$ $=\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}+XA,$ (34)
$g^{B}$ $=\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}-BX.$ (35)
In our method, we perform the standard forward and backward passes of LoRA, then adjust the gradients of A and B using Solutions (34) and (35), and subsequently update them. ∎
@@
Theorem.
When updating matrices $A$ and $B$ using the closed-form solution from Theorem 2.1, we proceed as follows: $A\leftarrow A-\gamma g^{A},$ (36) $B\leftarrow B-\gamma g^{B},$ (37) where $\gamma\geq 0$ denotes the learning rate. Our method ensures a decrease in the loss, akin to the standard gradient descent algorithm, expressed by: $\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}\leq 0$ (38)
Proof.
In summary, the proof of Theorem 2.2 is divided into two distinct parts. To begin with, we demonstrate that $\mathrm{d}L$ can be expressed in the following form:
$\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}.$ (39)
In the second part, we prove that this expression for $\mathrm{d}L$ is always less than or equal to zero: $\mathrm{d}L\leq 0$.
Part I. Therefore, in this part, we first prove Equation (39). During the optimization process, the differential change in the loss function, $\mathrm{d}L$, can be expressed in terms of the differentials $\mathrm{d}A$ and $\mathrm{d}B$ as follows:
$\mathrm{d}L=\langle\frac{\partial L}{\partial A},\mathrm{d}A\rangle_{F}+\langle\frac{\partial L}{\partial B},\mathrm{d}B\rangle_{F}.$ (40)
From Equation (36) and (37), we can derive that:
$\mathrm{d}A=-\gamma g^{A},\quad\mathrm{d}B=-\gamma g^{B}.$ (41)
Given that $\frac{\partial L}{\partial A}=g^{A}_{lora}$ and $\frac{\partial L}{\partial B}=g^{B}_{lora}$, it follows that:
$\mathrm{d}L$ $=-\gamma(\langle g^{A}_{lora},g^{A}\rangle_{F}+\langle g^{B}_{lora},g^{B}\rangle_{F})$ (42)
$=-\gamma(\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}$
$+\langle g^{A}_{lora},XA\rangle_{F}-\langle g^{B}_{lora},BX\rangle_{F}).$
And we have the following equation:
$\langle g^{A}_{lora},XA\rangle_{F}-\langle g^{B}_{lora},BX\rangle_{F}$ (43)
$=$ $\langle g^{A}_{lora}A^{T},X\rangle_{F}-\langle B^{T}g^{B}_{lora},X\rangle_{F}$
$=$ $\langle g^{A}_{lora}A^{T}-B^{T}g^{B}_{lora},X\rangle_{F}$
$=$ $\langle(sB^{T}g)A^{T}-B^{T}(sgA^{T}),X\rangle_{F}$
$=$ $0$
Therefore, we have:
$\mathrm{d}L=-\gamma\{\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}+\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\}.$ (44)
Part II. In this part, we aim to prove $\mathrm{d}L\leq 0$. Given that the learning rate $\gamma>0$, it suffices to show the following inequalities:
$\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}\geq 0,$ (45)
$\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\geq 0.$ (46)
By proving these inequalities, we can establish that $\mathrm{d}L\leq 0$ as derived from Equation (39).
① Proof of $\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}\geq 0$.
To begin with, we need to show that $(B^{T}B)^{-1}$ is positive definite. To establish this, it is sufficient to show that $B^{T}B$ is positive definite, as the inverse of a positive definite matrix is also positive definite. To achieve this, consider any non-zero vector $x$, and noting that $B$ is full-rank, we have,
$\langle x,B^{T}Bx\rangle=\langle Bx,Bx\rangle=\|Bx\|^{2}>0.$ (47)
This shows that $B^{T}B$ is positive definite. Consequently, $(B^{T}B)^{-1}$ is positive definite as well. Since $(B^{T}B)^{-1}$ is positive definite, and thus we can apply Cholesky decomposition, and $(B^{T}B)^{-1}=UU^{T}$. With this, we have,
$\langle g^{A}_{lora},\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}\rangle_{F}$ $=\frac{1}{s^{2}}\langle g^{A}_{lora},UU^{T}g^{A}_{lora}\rangle_{F}$ (48)
$=\frac{1}{s^{2}}\langle U^{T}g^{A}_{lora},U^{T}g^{A}_{lora}\rangle_{F}$
$=\frac{1}{s^{2}}\|U^{T}g^{A}_{lora}\|_{F}^{2}\geq 0$
② Proof of $\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}\geq 0$.
Similarly, we can prove that matrix $(AA^{T})^{-1}$ is positive-definite. By employing Cholesky decomposition, we express $(AA^{T})^{-1}=UU^{T}$, where $U$ is a lower-triangle matrix. Subsequently, we define $P=I-B(B^{T}B)^{-1}B^{T}$. It can be shown that $P^{2}=P$ and $P$ is symmetry, indicating that $P$ is a projection matrix. Consequently, the eigenvalues of $P$ are either 0 or 1, which implies that $P$ is positive semi-definite. Utilizing the Cholesky decomposition, we derive that $P=VV^{T}$, where $V$ is a lower-triangle matrix. Finally, we have:
$\langle g^{B}_{lora},\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}\rangle_{F}$ $=\frac{1}{s^{2}}\langle g^{B}_{lora},VV^{T}g^{B}_{lora}UU^{T}\rangle_{F}$ (49)
$=\frac{1}{s^{2}}\langle V^{T}g^{B}_{lora}U,V^{T}g^{B}_{lora}U\rangle_{F}$
$=\frac{1}{s^{2}}\|V^{T}g^{B}_{lora}U\|_{F}^{2}\geq 0$
In summary, based on the above proofs, we have demonstrated that:
[…]
B.4 Proof of Theorem 2.3
Theorem.
@@
Consider the optimization problem, $\min_{X}\|g^{A}-g^{A}_{lora}\|_{F}^{2}+\|g^{B}-g^{B}_{lora}\|_{F}^{2},$ (51) where $g^{A}$ and $g^{B}$ are the optimal solutions as stated in Theorem 2.1. The optimal $X$ can be determined by solving the Sylvester equation: $B^{T}BX+XAA^{T}=-\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}A^{T},$ (52) which has a unique solution $X$ provided that $B^{T}B$ and $-AA^{T}$ do not have any shared eigenvalues.
Proof.
For simplicity, we denote $L=\|g^{A}-g^{A}_{lora}\|_{F}^{2}+\|g^{B}-g^{B}_{lora}\|_{F}^{2}$. To solve the optimization problem, we need to satisfy the following conditions:
$\frac{\partial L}{\partial X}=0.$ (53)
Since $g^{A}$ and $g^{B}$ are solutions in Theorem 2.1 and $g^{A}_{lora}=sB^{T}g$ and $g^{B}_{lora}=sgA^{T}$, we obtain that:
$2(g^{A}-g^{A}_{lora})A^{T}$ $-2B^{T}(g^{B}-g^{B}_{lora})=0,$ (54)
$\Rightarrow\quad g^{A}A^{T}-B^{T}g^{B}$ $=g^{A}_{lora}A^{T}-B^{T}g^{B}_{lora},$
$\Rightarrow\quad B^{T}BX+XAA^{T}$ $=-\frac{1}{s^{2}}(B^{T}B)^{-1}g^{A}_{lora}A^{T},$
which is a Sylvester equation. This equation has a unique solution for $X$ if and only if $B^{T}B$ and $-AA^{T}$ have no shared eigenvalues.
@@ Appendix C Optimization Algorithms
In this section, we present the pseudo-codes for implementing our LoRA-Pro method using the SGD (Sutskever et al., 2013) and AdamW (Loshchilov & Hutter, 2019) optimizers. The details are provided in Algorithm 1, Algorithm 2, and Algorithm 3, respectively.
LoRA-Pro with SGD optimizer. In the standard SGD algorithm, as illustrated in Algorithm 1, all we need to do is adjusting the gradients of matrices $A$ and $B$ with the solutions in Theorem 2.1.
@@
LoRA-Pro with AdamW optimizer. In AdamW optimizer, the implementation becomes more complex. We offer two scalable implementation modes for LoRA-Pro with AdamW: efficient mode and full mode. We use the efficient mode by default. In efficient mode, the memory usage of LoRA-Pro’s optimizer remains the same as that of standard LoRA, but this mode tends to be less stable. In full mode, LoRA-Pro aligns more closely with full fine-tuning during the optimization process, resulting in higher memory usage but improved stability.
In efficient mode, we only need to modify the gradients of matrices $A$ and $B$. Then, we separately compute the first- and second-order gradient information for $A$ and $B$ before performing the update. The detailed algorithm is shown in Algorithm 2. In this mode, memory usage remains consistent with that of LoRA (Hu et al., 2022), delivering better results, though not as stable as full mode.
In full mode, we aim to closely approximate full fine-tuning during optimization. Several modifications are necessary. Firstly, in order to mimic full fine-tuning, after adjusting the gradients of matrices $A$ and $B$, we need to compute the equivalent gradient,
$\tilde{g}=sg^{B}A+sBg^{A}.$ (55)
$W\leftarrow(1-\gamma\lambda)(W_{0}+sBA).$ (57)
This can be decomposed into:
Ablation study on the AdamW optimizer. We conducted experiments to compare the memory cost differences between the two implementations of the AdamW optimizer, as well as their performance on MT-Bench, GSM8K, and HumanEval. The results of the study are presented in Table 7. As shown in the table, the full mode achieves better performance compared to the efficient mode. We believe this is because the full mode more closely aligns with full fine-tuning during the optimization process. However, as a trade-off, full mode incurs a higher GPU memory cost due to the calculation of the first- and second-order momentum information for the equivalent gradients $\tilde{g}$. Despite this, its memory usage is still significantly lower than that of full fine-tuning (which theoretically requires 56GB of memory). Therefore, we use efficient mode as the default setting for LoRA-Pro.
Table 7: Comparison of the AdamW optimizer implementations with LoRA-Pro. We present two types of AdamW optimizers for LoRA-Pro: the efficient mode, which maintains the same memory cost as the vanilla LoRA optimizer, and the full mode, which is more aligned with full fine-tuning and offers greater stability during training.
Memory Cost MT-Bench GSM8K HumanEval
Full FT $>$ 48 GB 5.30±0.11 59.36±0.85 35.31±2.13
LoRA-Pro (Efficient) 23.05 GB 5.86±0.06 54.23±0.79 22.76±0.35
LoRA-Pro (Full) 43.87 GB 6.05±0.25 54.06±0.46 23.98±0.35
Algorithm 2 LoRA-Pro with AdamW optimizer (Efficient Mode)
0:  Given initial learning rate $\gamma$, scaling factor $s$, and $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8},\lambda\in\mathbb{R}$
1:  Initialize time step $t\leftarrow 0$, low-rank matrices $A_{0}\in\mathbb{R}^{r\times n}$ and $B_{0}\in\mathbb{R}^{m\times r}$, first momentum $m^{A}_{0}\in\mathbb{R}^{r\times n}$, $m^{B}_{0}\in\mathbb{R}^{m\times r}$, second momentum $v^{A}_{t}\in\mathbb{R}^{r\times n}$, $v^{B}_{t}\in\mathbb{R}^{m\times r}$
2:  repeat
3:     $t\leftarrow t+1$
[…]
8:     $g^{B}=\frac{1}{s^{2}}[I-B(B^{T}B)^{-1}B^{T}]g^{B}_{lora}(AA^{T})^{-1}-BX$
9:     $m^{A}_{t}\leftarrow\beta_{1}m^{A}_{t-1}+(1-\beta_{1})g^{A}$
10:     $v^{A}_{t}\leftarrow\beta_{2}v^{A}_{t-1}+(1-\beta_{2})g^{A}*g^{A}$
11:     $\hat{m}^{A}_{t}\leftarrow\frac{m_{t}^{A}}{1-\beta_{1}^{t}}$
12:     $\hat{v}^{A}_{t}\leftarrow\frac{v_{t}^{A}}{1-\beta_{2}^{t}}$
13:     $\tilde{g}^{A}\leftarrow\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
14:     $A_{t}\leftarrow(1-\lambda\gamma)A_{t-1}$
15:     $A_{t}\leftarrow A_{t}-\gamma\tilde{g}^{A}$
16:     $m^{B}_{t}\leftarrow\beta_{1}m^{B}_{t-1}+(1-\beta_{1})g^{B}$
17:     $v^{B}_{t}\leftarrow\beta_{2}v^{B}_{t-1}+(1-\beta_{2})g^{B}*g^{B}$
18:     $\hat{m}^{B}_{t}\leftarrow\frac{m_{t}^{B}}{1-\beta_{1}^{t}}$
19:     $\hat{v}^{B}_{t}\leftarrow\frac{v_{t}^{B}}{1-\beta_{2}^{t}}$
20:     $\tilde{g}^{B}\leftarrow\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
21:     $B_{t}\leftarrow B_{t-1}-\gamma\tilde{g}^{B}$
22:  until stopping criterion is met
23:  return  optimized parameters $A_{t}$ and $B_{t}$
Algorithm 3 LoRA-Pro with AdamW optimizer (Full Mode)